A researcher claims that the average wind velocity that has to be considered in building the bridge in Batanes is 8 2miles per hour. A sample of 32 days has an average wind speed of 8.4 miles per hour different from the claim. The standard deviation of the population is 0.7 mile per hour. At a 01, 5 there enough evidence to support the claim? Use the P-value method.
Answer #1

Solution:
Null hypothesis H0: Mean = 8.2 miles per hour
Alternate hypothesis H1: Mean is not equal to 8.2 miles per hour
Given in the question

No. of sample = 32
sample mean = 8.4 miles per hour
Population standard deviation = 0.7 miles per hour
we will use normal distribution method so
Z = (8.4-8.2)/0.7/sqrt(32) = 0.2/0.7/5.66 = 1.313/0.7 = 1.62
from Z table we found p-value is 0.1052
as alpha = 0.05
we are failed to reject the null hypothesis as p-value is greater than alpha value. so we can say that building the bridge in bananas is not equal to 8.2 miles per hour.